OpenStudy (anonymous):

the length of the small leg of a 30-60-90 triangle is 5 what is the perimeter of the triangle? A.) 5+15√3 B.) 15 +5√3 C.)5/2 + 15/2 √3 D.) 15/2 + 5/2 √3

OpenStudy (anonymous):

a = 5. the sides of a 30-60-90 triangle are \[a, 2a, a \sqrt{3}\]

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

no can you help me understand more

OpenStudy (anonymous):

ok so you're finding perimeter, which is the sum of all the side lengths. to find the lengths just plug in the value of a (which is 5) into \[a, 2a, a \sqrt{3}\] and simplify

OpenStudy (anonymous):

\[5 + 2(5) + 5\sqrt{3}\]

OpenStudy (anonymous):

so is my answer B.)

OpenStudy (anonymous):

@turqoisetuna

OpenStudy (anonymous):

yes good job